# LED: LLM Enhanced Open-Vocabulary Object Detection without Human Curated Data Generation

### v1

@@ Abstract
Large foundation models trained on large-scale visual-text data can significantly enhance Open Vocabulary Object Detection (OVD) through data generation. However, this may lead to biased synthetic data and overfitting to specific configurations. It can sidestep biases of manually curated data generation by directly leveraging hidden states of Large Language Models (LLMs), which is surprisingly rarely explored. This paper presents a systematic method to enhance visual grounding by utilizing decoder layers of the LLM of a MLLM. We introduce a zero-initialized cross-attention adapter to enable efficient knowledge transfer from LLMs to object detectors, an new approach called L ED (LLM Enhanced Open-Vocabulary Object Detection). We demonstrate that intermediate hidden states from early LLM layers retain strong spatial-semantic correlations that are beneficial to grounding tasks. Experiments show that our adaptation strategy significantly enhances the performance on complex free-form text queries while remaining the same on plain categories. With our adaptation, Qwen2-0.5B with Swin-T as the vision encoder improves GroundingDINO by 2.33% on Omnilabel, at the overhead of 8.7% more GFLOPs. Qwen2-0.5B with a larger vision encoder can further boost the performance by 6.22%. We further validate our design by ablating on varied adapter architectures, sizes of LLMs, and which layers to add adaptation.
@@ 1 Introduction
Figure 1: Overview of open vocabulary object detection approaches. (a) Data generation methods prone to distribution shift and overfitting; (b) Our proposed adapter-based framework leverages hidden states from the MLLM to provide adaptation prompts to the detection decoder, enhancing semantic richness and generalization without relying on human curated data.
Open Vocabulary object Detection (OVD) localizes objects described by a free-form text, significantly increasing the size of the label space and enabling broader applications, compared to the traditional closed-set object detection with a fixed label space. Great prior efforts [44, 26, 15, 35, 45] focus on exploring large foundation models trained on image and/or text data to generate high-quality data for OVD, greatly boosting the performance on challenge benchmarks [10, 30, 33]. Their success demonstrates that large foundation models implicitly capture rich semantic knowledge and reasoning capabilities through pretraining, which are beneficial for accurately localizing objects under complex, free-form textual queries.
@@
In this paper, We present L ED , a new approach to leverage the knowledge of large foundation models for OVD without biased hand-crafted rules. Unlike prior studies that explicitly generate data under some assumptions, we believe the intermediate hidden states of large foundation models embody the essential knowledge and develop a novel framework to embed those hidden states into prevailing object detector. Specifically, as shown in Figure 1(a), existing approaches typically generate new data with specific goals, such as generating negative samples to address hallucination issues [37, 41, 29, 36, 46] or generating more detailed descriptions for spatial and logical relationships [15], which in turn limits the robustness and generalization of the model. In contrast as Figure 1(b), we extract hidden states from the LLM of a multimodal LLM (MLLM), which contains rich semantics and knowledge from large-scale pretraining, and integrate them into the object detector.
We apply L ED to popular open-source object detectors. GroundingDino [22]. Our method significantly improves the performance on prevailing benchmarks, including RefCOCO [39] and OmniLabel benchmark [30]. Besides, we find that hidden states from the first several layers of a MLLM are rich enough to improve the OVD models. Probably, MLLM focuses more on vision tokens in early layers and then on text tokens in later layers, as illustrated in section 3. Thus, vision tokens of deeper layers are not quite updated and embody similar knowledge as earlier ones. We believe our finding and analysis are beneficial for future studies that explore hidden states of large foundation models for downstream tasks.
Our contributions are summarized as. (1) We present a new framework that leverages knowledge from large foundation models. Instead of generating data, we use hidden states to improve OVD models. (2) We find and analyze that hidden states of the first several layers of a MLLM contains sufficient knowledge for detectors, which paves the path for future studies exploring hidden states for downstream tasks. (3) We conducted extensive experiments, and show that L ED can improve OVD models on various benchmarks and possesses significant zero-shot capabilities.
@@ 2 Relate Work
Multi-Modal Large Language Models: Recent advances in MLLMs have catalyzed research on visual knowledge integration, where modality alignment plays a pivotal role in vision-language learning [11]. CLIP pioneers this direction through contrastive learning on large-scale image-text pairs [27], while BLIP further improves alignment quality via sophisticated data curation strategies [14, 13]. LLaVA introduces visual instruction tuning using GPT-4-generated multimodal data, significantly advancing the general-purpose multimodal reasoning capabilities of large language-and-vision models[20]. Building upon these foundations, InternVL introduces a progressive alignment paradigm that effectively scales vision-language models, achieving state-of-the-art performance in both vision foundation models and VLLMs [4, 5, 6, 9, 32]. LLaMA Adapter introduces a lightweight adaptation method with a zero-initialized attention mechanism, enabling efficient instruction tuning of LLaMA using only 1.2M learnable parameters and achieving performance comparable to fully fine-tuned models [43]. It is noteworthy that LLaMA Adapter improves the performance of specific language tasks by providing the decoder with continuous feature representations of vocabulary as prompts. L ED aims to leverage the vision understanding of MLLM to directly provide high-level feature representation guidance for downstream tasks.
@@
However, synthetic data often diverges from real-world data distributions, introducing biases that impair model performance under natural variations and unseen conditions. Furthermore, generative techniques relying heavily on constructed datasets may inadvertently overfit specific dataset characteristics, restricting model adaptability when confronted with open-world scenarios or challenging distributions not adequately represented during training. Although methods like GLIP and DetCLIP [34] have enhanced concept grounding through enriched annotations and dictionaries, the intrinsic reliance on synthetic constructs remains a fundamental challenge.
@@ 3 Methodology
Our framework establishes an efficient pipeline for transferring multimodal knowledge from Large Language Models (LLMs) to OVD models via intermediate feature adaptation. As shown in Figure 2, we analyzed token attention changes across different layers of the LLM decoder for Qwen2-0.5B. In the early layers of the LLMs decoder (e.g., layers 0 to 4), the median attention score decreases significantly and fluctuates slightly in the subsequent 20 layers. This indicates that the MLLMs already exhibit strong attention or encoding capabilities for input image tokens in the early layers, while the later layers may focus more on fine-tuning task- or semantics-related text tokens. From a practical perspective, this distribution of attention scores suggests that early layers are critical for extracting vision-based features for downstream tasks such as object detection. Meanwhile, later layers can be selectively utilized or fused depending on specific task requirements.
@@
To validate this concept, we designed an architecture as shown in Figure 3. The architecture comprises four key components: (1) An efficient cross-modal alignment strategy in subsection 3.1; (2) A MLLM for semantic adaptation prompts generation in subsection 3.2; (3) A sufficiently universal end-to-end detector in subsection 3.3; (4) An adapter for feature fusion based on zero-initialized cross-attention in subsection 3.4.
@@ 3.4 Zero-initialized Cross-attention
A zero-initialized cross-modal attention mechanism is designed based on a small gating unit. It is worth noting that since the weights of the entire adapter are randomly initialized, directly providing the adaptation prompt to the detector decoder introduces disordered noise, which is detrimental to the training of the network. Given an decoder embedding $\mathbf{E_{D}}^{\ell_{D}}\in\mathbb{R}^{B\times T\times d}$ from $\ell_{D}$th decoder layer and adaptation pr ompts $\mathbf{A_{P}}\in\mathbb{R}^{B\times L\times d}$, our zero-initialized cross-modal attention mechanism operates as follows:
@@
$\tilde{\mathbf{S}}=$ $\mathcal{C}(\tanh(\mathbf{g})\odot\text{softmax}(\mathbf{S}_{:,0:L}),$
$\text{softmax}(\mathbf{S_{:,L+1:L+T}})\in\mathbb{R}^{B\times h\times T\times(L+T)}$
where RoPE is defined as $\tilde{\mathbf{Q}},\tilde{\mathbf{K}}=\text{RoPE}(\mathbf{Q},\mathbf{K};\Theta)$ with angular parameter matrix $\Theta$; $\mathbf{S}$ represents the raw attention score matrix, and $\tilde{\mathbf{S}}$ is attention weight matrix obtained by applying softmax; $\mathbf{g}\in\mathbb{R}^{1\times h\times 1\times 1}$ is a learnable gating vector initialized to $\mathbf{0}$. and an activation function $\tanh(\mathbf{\cdot})$ is adopted to regulate the scale of $\mathbf{g}$ to into $-1\sim 1$. If adaptation prompts are initialized randomly, they may introduce noise or interfere with word tokens during the initial stages of training, thereby compromising the stability and effectiveness of the fine-tuning process.
@@ 4 Experiments
In this section, we first detail datasets and implementation specifics, including MLLM feature alignment strategies and adapter training protocols (see subsection 4.1). Then, via evaluations on the OmniLabel benchmark[30], we analyze strategies for sharing vision decoder features (see subsubsection 4.2.1), investigate the impact of different adapter architectures (see subsubsection 4.2.2), and observe performance improvements across various scales of LLMs for visual detection and grounding tasks (see subsubsection 4.2.3). Finally, we conduct ablations to precisely quantify the contributions of MLLM decoder hidden states at various layers to detection performance (see subsection 4.3).
@@ 4.1 Dataset and Implementation Details
  1. 1.
    Stage 1: This phase leverages LAION-CC-SBU (558K samples) [18] to establish cross-modal alignment through diverse visual-linguistic interactions.
  2. 2.
    Stage 2: This phase combines ShareGPT4V (102K) and SFT (665K) [2] for instruction tuning, augmented with domain-specific datasets (ChartQA [18K] [24], AI2D [12K][12], DocVQA [10K] [25]) to enhance structured visual understanding.
Stage 3 - Adapter training: We employ Grounding DINO[21] pretrain on Object365, GoldG and Cap4M as a unified detection pre-training framework across three datasets: Objects365 [31], COCO2017, and Flickr30k [38]. We implement our framework using InternVL2-1B (Qwen2-0.5B backbone) as the MLLM [4, 32, 9, 5, 6] and Open-GroundingDINO [21] as the detection codebase.
Training is conducted on 8×A100 GPUs. We pretrain (Stage 1) and fine-tune (Stage 2) the MLP layer each for 1 epoch. For pretraining, the warmup ratio, batch size, learning rate, and weight decay are set to 0, 128, $1\times{10}^{-3}$, and 0.03, respectively. The comprehensive optimization process spanned roughly 4900 steps. For fine-tuning, the warmup ratio, batch size, learning rate, and weight decay are set to 0.03, 128, $4\times{10}^{-5}$, and 0, respectively. In Stage 3,We train the Adapter with a batch size of 64 for 1 epoch. We employ cosine learning rate scheduling (base lr=$2e^{-4}$ without warmup or weight decay. The MLP layer and LLMs is trained with the adapter using a different learning rate of $4\times{10}^{-5}$.
@@ 4.2 Evaluation
We evaluated and analyzed adapters under different architectures based on the OmniLabel benchmark[30]. Additionally, we compared the performance results of the hidden state acquisition scheme in InternVL2 with GroundingDINO (abbreviated as G-DINO in the table).
@@ 4.2.1 Vision Decoder Share from Detector
The rich semantic understanding in the MLLM primarily stems from the LLM rather than the vision encoder. Therefore, we first use the architecture and weights of the vision encoder with those from GroundingDINO to provide vision features to LLMs. We shared the architecture and weights of the vision encoder for both the detector and MLLM, and discarded a series of image preprocessing schemes commonly used in MLLM. Most MLLMs employ dynamic processing that dynamically selects an optimal aspect ratio from a set of predefined ratios. This process partitions the input image into pixel-based tiles and generates a thumbnail to retain global contextual information [19, 4, 32, 9, 5, 6]. But we discarded this to eliminate the additional image encoder and avoid providing extra fine-grained features through pre-segmented images, which could confuse the LLMs’ understanding of vision tokens. We first pretrain (stage 1 in Figure 3) and then finetune (stage 2 in Figure 3) the MLLM while freezing the vision encoder and LLM during training. subsection 6.4 (See supplement) provides an MME evaluation of the alignment degree of the MLP layers in stages 1 and 2.
The vision decoder we sharing from the detector is Swin-Tiny, introduce a pixel shuffle operation with a scaling factor of 4 to downsample the spatial dimensions and expand the channel dimension. Specifically, the height $h$ and width $w$ of the vision encoder’s output are reduced by a factor of 4, and the channel dimension is increased by a factor of 16. The output is then repeated and truncated along the channel dimension to match the input dimension of the MLP projection layer (4096). The alignment of feature spaces is achieved by pretraining the MLLM’s MLP projection layer as follows:
$\hat{f}_{v}=\text{MLP}(\text{Truncate}(\text{Repeat}$ (6)
$(\text{PixelShuffle}_{4}$ $(\text{Vision Encoder}(I)))))$
@@
Finally, when training the adapter (stage 3 in Figure 3), both the LLM and the MLP layers are trained simultaneously to adapt to downstream tasks. Image embeddings from Swin-Tiny (used in GroundingDINO) are fed into the LLM after alignment by a projector. The pretrained weights for Swin-Tiny, which are the same as those used in GroundingDINO and provided by Open-GroundingDINO [47], are utilized and frozen during every stages.
We measured the GFLOPs for each architecture and selected the $2_{th}$ decoder layer, removing subsequent layers. As shown in Table 1, L ED requires only 8.7% more FLOPs than GroundingDINO, with the majority of the computational cost attributed to the LLM’s decoder.
Model/Component Params (M) GFLOPS
G-DINO 172M 412G
Adapter 0.26M 1.47G
Conv Layer 2.06M 0.54G
Qwen2-0.5B $2_{th}$ 56M 34G
Total 58M 36G
Table 1: Model parameters and computational costs. It can be seen that almost all additional overhead comes from a lightweight (0.5B) LLM with only two layers, while the refinement of hidden states by the adapter is extremely efficient.
The OmniLabel benchmark [30] for our method is presented in Table 2 with an improvement of $2.33\%$ over GroungdingDINO. This improvement can be attributed to the adaptation prompts provided by the LLMs.
@@
 Swin-T + Qwen2-0.5B 24.01% $\uparrow$2.3333.01% $\downarrow$0.2518.53% $\uparrow$2.44%
Table 2: OmniLabel evaluation of the pipeline for stage 1,2,3 and share vision decoder from detector with improvements over GroundingDINO.

### v2

@@ Abstract
Large foundation models trained on large-scale vision–language data can boost Open-Vocabulary Object Detection (OVD) via synthetic training data, yet the hand-crafted pipelines often introduce bias and overfit to specific prompts. We sidestep this issue by directly fusing hidden states from Large Language Models (LLMs) into detectors—an avenue surprisingly under-explored. This paper presents a systematic method to enhance visual grounding by utilizing decoder layers of the LLM of a MLLM. We introduce a zero-initialized cross-attention adapter to enable efficient knowledge-fusion from LLMs to object detectors, an new approach called L ED (LLM Enhanced Open-Vocabulary Object Detection). We find that intermediate LLM layers already encode rich spatial semantics; adapting only the early layers yields most of the gain. With Swin-T as the vision encoder, Qwen2-0.5B + L ED lifts GroundingDINO by 3.82 % on OmniLabel at just 8.7 % extra GFLOPs, and a larger vision backbone pushes the improvement to 6.22 %. Extensive ablations on adapter variants, LLM scales and fusion depths further corroborate our design.
@@ 1 Introduction
Figure 1: Overview of open vocabulary object detection approaches. (a) Data generation methods prone to distribution shift and overfitting; (b) Our proposed adapter-based framework directly leverages knowledge from the MLLM to the detection decoder.
Open Vocabulary object Detection (OVD) localizes objects described by free-form text, significantly increasing the size of the label space and enabling broader applications, compared to the traditional closed-set object detection with a fixed label space. Significant prior efforts [43, 24, 14, 34, 44] focus on exploring large foundation models trained on image and/or text data to generate high-quality data for OVD, greatly boosting performance on challenge benchmarks [10, 28, 31]. Their success demonstrates that large foundation models implicitly capture rich semantic knowledge and reasoning capabilities through pretraining, which are beneficial for accurately localizing objects under complex, free-form textual queries.
@@
We argue that the latent representations of large multimodal language models already encode rich visual–linguistic concepts that need not be materialised into explicit labels. Building on this observation, we introduce L ED , an Latent Enhanced Open-Vocabulary Object Detection that realises a novel knowledge-fusion paradigm: instead of learning a detector from scratch or augmenting training with generated data, we directly fuse intermediate hidden states from a frozen MLLM into popular mainstream detector via a lightweight adapter. Specifically, as shown in Figure 1(a), existing approaches primarily distill LLM knowledge into the object detector via explicit data generation—such as synthesizing negative examples to mitigate hallucinations [36, 40, 27, 35, 45], or creating detailed annotations of spatial and logical relationships [14]. However, these explicit generation methods inherently constrain model robustness and generalizability. In contrast, as illustrated in Figure 1(b), our method directly fuses the high-dimensional semantic knowledge encapsulated within LLM hidden states into the detector, thereby preserving richer pretrained knowledge without explicit data augmentation.
We apply L ED  to popular open-source object detectors. GroundingDino [19]. Our method significantly improves performance on prevailing benchmarks, including RefCOCO [38] and OmniLabel benchmark [28]. Besides, we find that hidden states from the first several layers of a MLLM are rich enough to improve the OVD models. Probably, MLLMs focuses more on vision tokens in early layers and then on text tokens in later layers, as illustrated in Section 3. Thus, vision tokens of deeper layers are not quite updated and embody similar knowledge as earlier ones.
Our contributions are summarized as:
(1) We introduce a lightweight adapter that directly fuses frozen MLLM hidden representations into any object detector, establishing a unified knowledge-injection pipeline and eliminating the need for expensive data synthesis.
(2) We systematically explore various fusion designs—layer selection, injection modality, adapter structure, etc.—and identify the most effective pipeline, offering clear guidelines for future work.
(3) We conduct extensive experiments, demonstrating that our approach not only improves OVD models across various benchmarks, but also achieves benefits comparable or even better to those obtained through data generation.
@@ 2 Relate Work
Multi-Modal Large Language Models: Recent MLLMs have driven visual–language integration through modality alignment. CLIP pioneered contrastive learning on large-scale image–text pairs [25], and BLIP enhanced alignment with curated data [12, 13]. LLaVA applies GPT-4–generated multimodal instruction tuning to boost reasoning [18]. InternVL’s progressive alignment scales vision–language models to SOTA in foundation models [4, 5, 6, 9, 30]. LLaMA Adapter uses a zero-initialized attention mechanism with just 1.2 M parameters for prompt-based tuning, matching full fine-tuning [42]. L ED  instead fuses high-level MLLM embeddings directly into downstream detectors.
@@
However, synthetic data often misaligns with real-world distributions, introducing biases that degrade performance under natural variations and unseen scenarios. Moreover, overreliance on generated samples can lead to overfitting to specific dataset artifacts, limiting adaptability in truly open-world settings. While approaches like GLIP [15] and DetCLIP [33] improve grounding via enriched annotations, their dependence on synthetic constructs remains a core limitation.
@@ 3 Methodology
Our framework establishes an efficient pipeline for transferring multimodal knowledge from LLMs to OVD models via intermediate feature adaptation. As shown in Figure 2, we analyze token attention changes across different layers of the LLM decoder for Qwen2-0.5B. In the early layers of the LLMs decoder (e.g., layers 0 to 4), the median attention score decreases significantly and fluctuates slightly in the subsequent 20 layers. This indicates that the MLLMs already exhibit strong attention or encoding capabilities for input image tokens in the early layers, while the later layers may focus more on fine-tuning task- or semantics-related text tokens Section A.1 (in the supplement).
To validate this concept, we design an architecture illustrated in Figure 3. The architecture comprises four key components: (1) an efficient cross-modal alignment strategy in Section 3.1; (2) an MLLM for semantic adaptation prompts generation in Section 3.2; (3) a universal end-to-end detector in Section 3.3; (4) a zero-initialized cross-attention adapter for feature fusion Section 3.4.
@@ 3.4 Zero-initialized Cross-attention
A zero-initialized cross-modal attention mechanism is designed based on a small gating unit. It is worth noting that since the weights of the entire adapter are randomly initialized, directly providing the adaptation prompt to the detector decoder introduces disordered noise, which is detrimental to the training of the network. Given a decoder embedding $\mathbf{E_{D}}^{\ell_{D}}\in\mathbb{R}^{B\times T\times d}$ from $\ell_{D}$th decoder layer and adaptation prompts $\mathbf{A_{P}}\in\mathbb{R}^{B\times L\times d}$, our zero-initialized cross-modal attention mechanism operates as follows:
@@
$\tilde{\mathbf{S}}=$ $\mathcal{C}(\tanh(\mathbf{g})\odot\text{softmax}(\mathbf{S}_{:,0:L}),\text{softmax}(\mathbf{S_{:,L+1:L+T}}))\in\mathbb{R}^{B\times h\times T\times(L+T)}$
where $\tilde{\mathbf{Q}},\tilde{\mathbf{K}}$ is defined as $\text{RoPE}(\mathbf{Q},\mathbf{K};\Theta)$ with angular parameter matrix $\Theta$; $\mathbf{S}$ represents the raw attention score matrix, and $\tilde{\mathbf{S}}$ is attention weight matrix obtained by applying softmax; $\mathbf{g}\in\mathbb{R}^{1\times h\times 1\times 1}$ is a learnable gating vector initialized to $\mathbf{0}$. and an activation function $\tanh(\mathbf{\cdot})$ is adopted to regulate the scale of $\mathbf{g}$ to into $-1\sim 1$. If adaptation prompts are initialized randomly, they may introduce noise or interfere with word tokens during the initial stages of training, thereby compromising the stability and effectiveness of the fine-tuning process.
@@ 4 Experiments
In this section, we first detail the datasets and implementation specifics Section 4.1). Next, through evaluations on the OmniLabel benchmark[28], we analyze strategies for sharing vision decoder features (see Section 4.2.1), investigate the impact of different adapter architectures (see Section 4.2.3), and observe performance improvements across various scales of LLMs for visual detection and grounding tasks (see Section 4.2.3). Finally, we conduct ablations to precisely quantify the contributions of MLLM decoder hidden states at various layers to detection performance (see Section 4.4).
@@ 4.1 Dataset and Implementation Details
  1. Stage 1: We leverages LAION-CC-SBU (558K samples) [17] to establish cross-modal alignment through diverse visual-linguistic interactions.
  2. Stage 2: We combine ShareGPT4V (102K) and SFT (665K) [3] for instruction tuning, augmented with domain-specific datasets Section D.1 (see Supplement) to enhance structured visual understanding.
Stage 3 - Adapter training: We employ Grounding DINO [19] pretrain on Object365, GoldG and Cap4M as a unified detection pre-training framework across three datasets: Objects365 [29], COCO2017, and Flickr30k [37]. We implement our framework using InternVL2-1B (Qwen2-0.5B backbone) as the MLLM [4, 30, 9, 5, 6] and Open-GroundingDINO [19] as the detection codebase. ALL experiment configuration As shown in Section A.3 (see Supplement)
@@ 4.2 Evaluation
We evaluated and analyzed adapters under different architectures based on the OmniLabel benchmark [28]. Additionally, we compared the performance results of the hidden state acquisition scheme in InternVL2 with GroundingDINO (abbreviated as G-DINO in tables).
@@ 4.2.1 Vision Decoder Share from Detector
The rich semantic understanding of the MLLM primarily originates from the LLM rather than the vision encoder. Thus, we adopt GroundingDINO’s vision encoder architecture and weights to provide visual inputs for both the detector and MLLM. We conduct MLLM training in two stages (see Figure 3): pretraining (stage 1) and finetuning (stage 2), freezing both the vision encoder and LLM during training. Section A.5 (supplementary) provides MME evaluations of alignment progress between stages.
Specifically, the shared Swin-Tiny vision encoder output undergoes a pixel shuffle operation (scaling factor 4), reducing spatial dimensions $(h,w)$ by a factor of 4 while expanding the channels by a factor of 16. This transformed output is then truncated to match the MLP projection layer dimension (4096), aligning the feature spaces through pretrained MLP layers:
$\hat{f}_{v}=\text{MLP}(\text{Truncate}(\text{Repeat}(\text{PixelShuffle}_{4}(\text{Vision Encoder}(I)))))$ (6)
@@
Table 1: Model params, compute cost, and inference latency; most overhead stems from the LLM’s two layers, with minimal adapter cost.
Framework Params GFLOPs Latency
G-DINO 172M 412G 299.3 ms
L ED +58M +36G +52.3ms
  -Adapter +2.2M +2.0G +2.0ms
  -Qwen2-0.5B $2_{L}$ +56M +34G +50.5ms
Table 2: OmniLabel evaluation for the three-stage pipeline and shared vision decoder, relative to GroundingDINO.
@@
Swin-T+ Qwen2-0.5B 25.51% $\uparrow$3.8233.01% $\uparrow$0.2520.28% $\uparrow$4.19%
Finally, when training the adapter (stage 3 in Figure 3), both the LLM and the MLP layers are trained simultaneously to adapt to downstream tasks. Image embeddings from Swin-Tiny (used in GroundingDINO) are fed into the LLM after alignment by a projector. The pretrained weights for Swin-Tiny, which are the same as those used in GroundingDINO and provided by Open-GroundingDINO [46], are utilized and frozen during every stage.
We measured the GFLOPs for each architecture and selected the $2_{nd}$ decoder layer, removing subsequent layers. As shown in Table 1, L ED  requires only 8.7% more FLOPs than GroundingDINO, with the majority of the computational cost attributed to the LLM’s decoder.
The OmniLabel benchmark [28] for our method is presented in Table 2 with an improvement of $3.82\%$ over GroundingDINO. This gain stems from the adaptation prompts generated by the LLM.